what are conjugated alkenes?

Grade:12

2 Answers

bhanuveer danduboyina
12 years ago

Conjugated alkenes are those where there are multiple double bonds which are on alternating carbons on a continuous chain. In non conjugated alkenes, any multiple double bonds are separated by at least one CH2 group. For instance, CH3CH=CH-CH=CH2 is a conjugated pentadiene, while CH2=CH-CH2-CH=CH2 is a non conjugated pentadiene. This has drastic effects on the reactivity of the alkenes (especially in free radical reactions) as well as spectroscopic and color effects, in extremely conjugated molecules.

PLEASE APPROVE MY ANSWER IF YOU LIKED IT!!

Godfrey Classic Prince
12 years ago

Dear sarthak mitra,

Conjugated alkenes are alkenes that have alternate single & double bonds between the carbon atoms of the molecule in its entire structure .

For example, 2-Butene has alternate single and double bonds between the carbon atoms in it .

 

This is an example of conjugated alkenes .

Hope this helped you immensely...

All the Very Best & Good Luck to you ....

Regards,

AskIITians Expert,

Godfrey Classic Prince

IIT-Madras

Please approve my answer if you liked it by clicking on "Yes" given below....!!Smile